OpenStudy (anonymous):

Solve. 9sqrt(x+3)-5=1 1st step: add 5 to both sides right? 9sqrt(x+3)= 6 2nd step: lost....

OpenStudy (saifoo.khan):

get rid of 9.

OpenStudy (saifoo.khan):

sqrt(x+3) = 6/9 sqrt(x+3) = 2/3 Take square on both sides to get rid of square root now. x+3 = (2/3)^2 Solve. (:

OpenStudy (anonymous):

darn fractions. throwin me off

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

would (2/3)^2 be simplified as 4/9?

OpenStudy (saifoo.khan):

Wait! Yup! I missed the denominator. lol. \[\Huge \color{red} \checkmark\]

OpenStudy (jiteshmeghwal9):